OpenStudy (anonymous):

Jan is five years older than Leann. Three years ago, the sum of their ages was 35. Find the age of each. How do I solve this?

OpenStudy (anonymous):

J=L+5 (Jan five years older than Leann) (J-3)+(L-3)=35 (3 years ago the sum of their ages was 35) J-L=5 J+L=41 add the equations 2J=46 J=23 L=18

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem :)